# Provably Efficient Multi-Objective Bandit Algorithms under Preference-Centric Customization

### v1

@@ Appendix D Analyses for Section 6 (Hidden Preference)
Our main result of Theorem 6 in Section 6 indicates that the proposed PUCB-HPM under hidden preference environment achieves sublinear expected regret $R(T)\leq\tilde{\mathcal{O}}(D\sqrt{T})$. To prove this, we need two key components. The first is to show that the value of $\hat{\boldsymbol{r}}_{i,t}$, the matrix of $\boldsymbol{V}_{t}$, and the region of $\Theta_{t}$ are good estimators of $\boldsymbol{\mu}_{i}$, $\mathbb{E}[\boldsymbol{V}_{t}]$ and $\overline{\boldsymbol{c}}$ respectively. The second is to show that as long as the aforementioned high-probability event holds, we have some control on the growth of the regret. We show the analyses regarding these two components in the following sections.
@@ D.1 Proof of Lemma 3
By Assumption 6.2 and Lemma 11, we have that that for any $d\in[D]$ and any $t$, $\boldsymbol{\zeta}_{t}(d)=\boldsymbol{c}_{t}-\boldsymbol{\overline{c}}$ is also $R$ sub-Gaussian: $\mathbb{E}[e^{x\boldsymbol{\zeta}_{t}(d)}]\leq e^{\frac{x^{2}R^{2}}{2}}$. For $\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$, the independence implies:
@@ D.2 Proof of Lemma 4 (Confidence Ellipsoid for $\boldsymbol{\overline{c}}$)
First we state two lemmas from [19] that will be utilized in our confidence analysis of preference estimator:
@@
Lemma 12 (Self-Normalized Bound for Vector-Valued Martingales [19], Theorem 1).
@@
Let $\{\mathcal{F}_{t}\}_{t=0}^{\infty}$ be a filtration, and let $\{\zeta_{t}\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\zeta_{t}$ is $\mathcal{F}_{t}$-measurable, $\mathbb{E}[\zeta_{t}\mid\mathcal{F}_{t-1}]$ = 0 and $\zeta_{t}$ is conditionally $R$-sub-Gaussian for some $R\geq 0$. Let $\{\boldsymbol{X}_{t}\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $\boldsymbol{X}_{t}$ is $\mathcal{F}_{t-1}$-measurable. Assume that $\boldsymbol{V}\in\mathbb{R}^{d\times d}$ is a positive definite matrix, and define $\boldsymbol{\overline{V}}_{t}=\boldsymbol{V}+\sum_{\ell=1}^{t}\boldsymbol{X}_{\ell}\boldsymbol{X}_{\ell}^{\top}$. Then for any $\alpha\geq 0$, with probability at least $1-\alpha$, for all $t\geq 1$, we have
@@
$\|\sum_{\ell=1}^{t}\zeta_{\ell}\boldsymbol{X}_{\ell}\|_{\boldsymbol{\overline{V}}_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\det\left(\boldsymbol{\overline{V}}_{t}\right)^{\frac{1}{2}}\det\left(\boldsymbol{V}\right)^{-\frac{1}{2}}}{\alpha}\right).$
Lemma 13 (Determinant-Trace Inequality [19], Lemma 10).
@@
Suppose $\boldsymbol{X}_{1},...,\boldsymbol{X}_{t}\in\mathbb{R}^{d}$ and $\|\boldsymbol{X}_{\ell}\|_{2}\leq L,\forall\ell\in[1,t]$. Let $\boldsymbol{\overline{V}}_{t}=\lambda\boldsymbol{I}+\sum_{\ell=1}^{t}\boldsymbol{X}_{\ell}\boldsymbol{X}_{\ell}^{\top}$ for some $\lambda>0$, then
@@
[…]
@@ D.3 Proof of Lemma 5 (Upper Confidence Bound for Expected Overall Reward)
(with probability at least $\qquad\qquad\qquad(\text{with probability at least }1-\vartheta)$
@@ D.4 Uniform Confidence Bound for Estimators
[…]
@@
$\mathbb{P}(B^{\mathsf{c}})$ =($=\mathbb{P}\left(\exists\{i,d,t\},|\hat{\boldsymbol{r}}_{i,t}(d)-\boldsymbol{\mu}_{i}(d)|>\sqrt{\frac{2\log(t/\alpha)}{N_{i,t}}}\right)$ (44)
$\leq 2\sum_{t=1}^{T}\sum_{i=1}^{K}\sum_{d=1}^{D}\mathbb{P}\left(|\hat{\boldsymbol{r}}_{i,t}(d)-\boldsymbol{\mu}_{i}(d)|>\sqrt{\frac{\log(t/\alpha)}{N_{i,t}}}\right)$
@@
$\mathbb{P}(C^{\mathsf{c}})$ =($=\mathbb{P}\left(\exists\{i,t,m,n\},\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t-1}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right](m,n)-\sum_{\ell\in\mathcal{T}_{i,t-1}}\left(w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right)(m,n)>\omega\sqrt{N_{i,t}\log\left(\frac{t}{\alpha}\right)}\right)$ (45)
$\leq\sum_{t=1}^{T}\sum_{i=1}^{K}\sum_{m=1}^{D}\sum_{n=m}^{D}\mathbb{P}\left(\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t-1}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right](m,n)-\sum_{\ell\in\mathcal{T}_{i,t-1}}\left(w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right)(m,n)>\omega\sqrt{N_{i,t}\log\left(\frac{t}{\alpha}\right)}\right)$
@@ D.5 Proof of Theorem 6 (Regret Analysis of Algorithm 2)
$R(T)=\sum_{t=1}^{T}\text{regret}_{t}\leq M+\sum_{t=M+1}^{T}\text{regret}_{t},$ (46)
where $\text{regret}_{t}$ denotes the instantaneous regret of PRUCB-HPM at step $t\in\left[T\right]$, and the last inequality holds since the fact that the instantaneous regret is upper-bounded by 1 (by Assumption 6.1).
@@
$R(T)$ $\leq M+\sum_{t=M+1}^{T}\text{regret}_{t}$ (48)
$\leq M+\sum_{t=M+1}^{T}\Big{(}\text{regret}_{t}^{\tilde{\boldsymbol{c}}}+\text{regret}_{t}^{\tilde{\boldsymbol{r}}}\Big{)}$
$\leq M+\underbrace{\sum_{t=M+1}^{T}\min\left(2B_{a_{t},t}^{c},1\right)}_{R_{M+1:T}^{\tilde{\boldsymbol{c}}}}+\underbrace{\sum_{t=M+1}^{T}2\|\boldsymbol{\overline{c}}\|_{1}\sqrt{\frac{\log(t/\alpha)}{N_{i,t}}}}_{R_{M+1:T}^{\tilde{\boldsymbol{r}}}},$
@@
$\sum_{t=M+1}^{T}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)\leq\sqrt{\frac{D}{2\log(9/8)}(T-M)\log\Big{(}1+\frac{\omega}{\lambda}(T-M)\Big{)}}.$
@@
$\sum_{t=M+1}^{T}\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{a_{t},t}}}\leq 2\sqrt{K(T-M)\log\left(\frac{T}{\alpha}\right)}.$
@@
$R_{M+1:T}^{\tilde{\boldsymbol{c}}}$ $\leq\sum_{t=M+1}^{T}\text{reg}_{t}^{(1)}+\sum_{t=M+1}^{T}\text{reg}_{t}^{(2)}$ (51)
$\leq 2\sqrt{2}\beta_{T}\sum_{t=M+1}^{T}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)+\frac{2\beta_{T}}{\sqrt{\lambda}}\sum_{t=M+1}^{T}\rho_{a_{t},t}$
@@
$R_{M+1:T}^{\tilde{\boldsymbol{r}}}=2\sum_{t=M+1}^{T}\|\boldsymbol{\overline{c}}\|_{1}\sqrt{\frac{\log(t/\alpha)}{N_{i,t}}}\underset{(a)}{\leq}2\sum_{t=M+1}^{T}\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{a_{t},t}}}\underset{(b)}{\leq}4\sqrt{K(T-M)\log\left(\frac{T}{\alpha}\right)},$ (52)
@@
This concludes the proof of Theorem 6. ∎

### v2

@@ Appendix D Analyses for Section 4 (Hidden Preference)
Our main result of Theorem 1 in Section 4 indicates that the proposed PUCB-HPM under hidden preference environment achieves sublinear expected regret $R(T)\leq\tilde{\mathcal{O}}(D\sqrt{T})$. To prove this, we need two key components. The first is to show that the value of $\hat{\boldsymbol{r}}_{i,t}$, the matrix of $\boldsymbol{V}_{t}$, and the value of $\boldsymbol{\hat{c}}_{t}$ are good estimators of $\boldsymbol{\mu}_{i}$, $\mathbb{E}[\boldsymbol{V}_{t}]$ and $\overline{\boldsymbol{c}}$ respectively. The second is to show that as long as the aforementioned high-probability event holds, we have some control on the growth of the regret. We show the analyses regarding these two components in the following sections. In the following proofs, we focus on the user-specific regret $R^{n}(T)$ and then sum the result over whole user set $N$. For notational simplicity, we fix a user $n$ and omit the user index $n$ as superscript.
@@ D.1 Proof of Lemma 1
By Assumption 5 and Lemma 5, we have that that for any $d\in[D]$ and any $t$, $\boldsymbol{\zeta}_{t}(d)=\boldsymbol{c}_{t}-\boldsymbol{\overline{c}}$ is also $R$ sub-Gaussian: $\mathbb{E}[e^{x\boldsymbol{\zeta}_{t}(d)}]\leq e^{\frac{x^{2}R^{2}}{2}}$. For $\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$, the independence implies:
@@ D.2 Proof of Lemma 2 (Confidence Ellipsoid for $\boldsymbol{\overline{c}}$)
First we state two useful lemmas that will be utilized in our confidence analysis of preference estimator:
@@
Lemma 6 (Self-Normalized Bound for Vector-Valued Martingales, Variant of Theorem 1 in (Abbasi-Yadkori, Pál, and Szepesvári 2011)).
@@
Let $\{\mathcal{F}_{t}\}_{t=0}^{\infty}$ be a filtration, and let $\{\zeta_{t}^{\prime}=\frac{\sqrt{\omega}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}}{\|\boldsymbol{r}_{a_{t},t}\|_{2}}\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\zeta_{t}$ is $\mathcal{F}_{t}$-measurable, $\mathbb{E}[\boldsymbol{\zeta_{t}}\mid\mathcal{F}_{t-1}]=\boldsymbol{0}$ and $\boldsymbol{\zeta_{t}}$ is conditionally $R$-sub-Gaussian for some $R\geq 0$. Let $\{\boldsymbol{r}_{t}^{\prime}=\frac{\sqrt{\omega}\boldsymbol{r}_{a_{t},t}}{\|\boldsymbol{r}_{a_{t},t}\|_{2}}\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $\boldsymbol{r}_{t}$ is $\mathcal{F}_{t}$-measurable. Assume that $\boldsymbol{V}\in\mathbb{R}^{d\times d}$ is a positive definite matrix, and define $\boldsymbol{\overline{V}}_{t}=\boldsymbol{V}+\sum_{\ell=1}^{\top}\boldsymbol{r}_{a_{\ell},\ell}^{\prime}{\boldsymbol{r}_{a_{\ell},\ell}^{\prime}}^{\top}$. Then for any $\alpha\geq 0$, with probability at least $1-\alpha$, for all $t\geq 1$, we have
@@
$\left\|\sum_{\ell=1}^{\top}\zeta_{\ell}^{\prime}\boldsymbol{r}_{a_{\ell},\ell}\right\|_{\boldsymbol{\overline{V}}_{t}^{-1}}^{2}\leq 2\omega R^{2}\log\left(\frac{\det\left(\boldsymbol{\overline{V}}_{t}\right)^{\frac{1}{2}}\det\left(\boldsymbol{V}\right)^{-\frac{1}{2}}}{\alpha}\right).$
Proof.
By definition and Tower property (Lemma 18), for any $t\in[T]$, we have
$\mathbb{E}[\zeta_{t}^{\prime}\boldsymbol{r_{a_{t},t}^{\prime}}|\mathcal{F}_{t-1}]=\mathbb{E}\left[\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}\mid\mathcal{F}_{t-1}\right]=\mathbb{E}\left[\mathbb{E}\left[\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}\mid\boldsymbol{r}_{a_{t},t},\mathcal{F}_{t-1}\right]\mid\mathcal{F}_{t-1}\right].$
By Assumption 3 (independence between reward and preference), we have $(\boldsymbol{r}_{a_{t},t}\perp\boldsymbol{\zeta}_{t})\mid\mathcal{F}_{t-1}$, and since $\boldsymbol{r}_{a_{t},t}$ is known inside the inner expectation, we have
$\mathbb{E}[\boldsymbol{\zeta}_{t}\mid r_{a_{t},t},\mathcal{F}_{t-1}]=\mathbb{E}[\boldsymbol{\zeta}_{t}\mid\mathcal{F}_{t-1}]=\boldsymbol{0}.$
Hence
$\mathbb{E}\left[\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}\mid\boldsymbol{r}_{a_{t},t},\mathcal{F}_{t-1}\right]=\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}\cdot\mathbb{E}[\boldsymbol{\zeta}_{t}\mid\boldsymbol{r}_{a_{t},t},\mathcal{F}_{t-1}]=\boldsymbol{0}$
which implies $\mathbb{E}[\boldsymbol{\zeta}_{t}^{\prime}\boldsymbol{r}_{a_{t},t}^{\prime}\mid\mathcal{F}_{t-1}]=\boldsymbol{0}$. Thus, the process $S_{t}=\sum_{\ell=1}^{\top}\boldsymbol{\zeta}_{\ell}^{\prime}\boldsymbol{r}_{a_{\ell},\ell}^{\prime}$ remains a martingale with respect to $\{\mathcal{F}_{t}\}$, even though $\boldsymbol{r}_{a_{\ell},\ell}$ is not $\mathcal{F}_{\ell-1}$-measurable.
Please note this differs from (Abbasi-Yadkori, Pál, and Szepesvári 2011), where the feature vector $X_{t}$ must be $\mathcal{F}_{t-1}$measureable so that $\mathbb{E}[\eta_{t}\boldsymbol{X}_{t}\mid\mathcal{F}_{t-1}]=\boldsymbol{X}_{t}\mathbb{E}[\eta_{t}\mid\mathcal{F}_{t-1}]=\boldsymbol{0}$, with $\eta_{t}$ as conditionally zero-mean sub-gaussian random variable.
With the conditions that $S_{t}$ is martingale and $\boldsymbol{\zeta}_{t}$ is conditionally sub-Gaussian, the process $M_{t}^{u}=\exp(\boldsymbol{u}^{\top}S_{t}-\frac{1}{2}\|\boldsymbol{u}\|_{V_{t}}^{2})$ is a nonnegative supermartingale w.r.t $\{F_{t}\}$ for arbitrary $\boldsymbol{u}$. Hence, Lemma 8 and Lemma 9 in (Abbasi-Yadkori, Pál, and Szepesvári 2011) can be directly applied to derive the self-normalized bound stated as Lemma 6 in our paper. The proof is omitted here, as it follows identically from the proof presented in Lemma 9 of Abbasi-Yadkori, Pál, and Szepesvári (2011). ∎
Lemma 7 (Determinant-Trace Inequality (Abbasi-Yadkori, Pál, and Szepesvári 2011), Lemma 10).
@@
Suppose $\boldsymbol{X}_{1},...,\boldsymbol{X}_{t}\in\mathbb{R}^{d}$ and $\|\boldsymbol{X}_{\ell}\|_{2}\leq L,\forall\ell\in[1,t]$. Let $\boldsymbol{\overline{V}}_{t}=\lambda\boldsymbol{I}+\sum_{\ell=1}^{\top}\boldsymbol{X}_{\ell}\boldsymbol{X}_{\ell}^{\top}$ for some $\lambda>0$, then
@@
$\qquad\qquad\text{(with probability at least }1-\alpha)$
@@ D.3 Proof of Lemma 3 (Upper Confidence Bound for Expected Overall Reward)
$\qquad\qquad\qquad(\text{with probability at least }1-\vartheta)$
@@ D.4 Uniform Confidence Bound for Estimators
$\text{Event A:}\Bigg\{|\boldsymbol{\mu}_{i}(d)-\boldsymbol{\hat{r}}_{i,t}(d)|\leq\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{i,t}}},\forall i\in[K],\forall d\in[D]\Bigg\},$
$\text{Event B:}\Bigg\{\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t-1}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right](m,n)-\sum_{\ell\in\mathcal{T}_{i,t-1}}\left(w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right)(m,n)\leq\omega\sqrt{N_{i,t}\log\left(\frac{t}{\alpha}\right)},$
@@
$\mathbb{P}(B^{\mathsf{c}})$ $=\mathbb{P}\left(\exists\{i,d,t\},|\hat{\boldsymbol{r}}_{i,t}(d)-\boldsymbol{\mu}_{i}(d)|>\sqrt{\frac{2\log(t/\alpha)}{N_{i,t}}}\right)$ (21)
$\leq 2\sum_{t=1}^{\top}\sum_{i=1}^{K}\sum_{d=1}^{D}\mathbb{P}\left(|\hat{\boldsymbol{r}}_{i,t}(d)-\boldsymbol{\mu}_{i}(d)|>\sqrt{\frac{\log(t/\alpha)}{N_{i,t}}}\right)$
@@
$\mathbb{P}(C^{\mathsf{c}})$ $=\mathbb{P}\left(\exists\{i,t,m,n\},\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t-1}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right](m,n)-\sum_{\ell\in\mathcal{T}_{i,t-1}}\left(w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right)(m,n)>\omega\sqrt{N_{i,t}\log\left(\frac{t}{\alpha}\right)}\right)$ (22)
$\leq\sum_{t=1}^{\top}\sum_{i=1}^{K}\sum_{m=1}^{D}\sum_{n=m}^{D}\mathbb{P}\left(\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t-1}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right](m,n)-\sum_{\ell\in\mathcal{T}_{i,t-1}}\left(w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right)(m,n)>\omega\sqrt{N_{i,t}\log\left(\frac{t}{\alpha}\right)}\right)$
@@ D.5 Proof of Theorem 1 (Regret Analysis of Algorithm 1)
$R(T)=\sum_{t=1}^{\top}\text{regret}_{t}\leq M+\sum_{t=M+1}^{\top}\text{regret}_{t},$ (23)
where $\text{regret}_{t}$ denotes the instantaneous regret of PRUCB-HPM at step $t\in\left[T\right]$, and the last inequality holds since the fact that the instantaneous regret is upper-bounded by 1 (by Assumption 4).
@@
$R(T)$ $\leq M+\sum_{t=M+1}^{\top}\text{regret}_{t}$ (25)
[…]
$\leq M+\underbrace{\sum_{t=M+1}^{\top}\min\left(2B_{a_{t},t}^{c},1\right)}_{R_{M+1:T}^{\tilde{\boldsymbol{c}}}}+\underbrace{\sum_{t=M+1}^{\top}2\|\boldsymbol{\overline{c}}\|_{1}\sqrt{\frac{\log(t/\alpha)}{N_{i,t}}}}_{R_{M+1:T}^{\tilde{\boldsymbol{r}}}}$,
@@
$\sum_{t=M+1}^{\top}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)\leq\sqrt{\frac{D}{2\log(9/8)}(T-M)\log\Big(1+\frac{\omega}{\lambda}(T-M)\Big)}.$
@@
$\sum_{t=M+1}^{\top}\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{a_{t},t}}}\leq 2\sqrt{K(T-M)\log\left(\frac{T}{\alpha}\right)}.$
@@
$R_{M+1:T}^{\tilde{\boldsymbol{c}}}$ $\leq\sum_{t=M+1}^{\top}\text{reg}_{t}^{(1)}+\sum_{t=M+1}^{\top}\text{reg}_{t}^{(2)}$ (28)
$\leq 2\sqrt{2}\beta_{T}\sum_{t=M+1}^{\top}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)+\frac{2\beta_{T}}{\sqrt{\lambda}}\sum_{t=M+1}^{\top}\rho_{a_{t},t}$
@@
$R_{M+1:T}^{\tilde{\boldsymbol{r}}}=2\sum_{t=M+1}^{\top}\|\boldsymbol{\overline{c}}\|_{1}\sqrt{\frac{\log(t/\alpha)}{N_{i,t}}}\underset{(a)}{\leq}2\sum_{t=M+1}^{\top}\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{a_{t},t}}}\underset{(b)}{\leq}4\sqrt{K(T-M)\log\left(\frac{T}{\alpha}\right)},$ (29)
@@
Summing above result over user set $[N$ concludes the proof of Theorem 1. ∎